OpenStudy (anonymous):

A 5.0nC point charge travels in a circle of radius 1.0cm. What is its speed so that the current it generates is 1.0mA?

OpenStudy (anonymous):

Let the speed be V. hence time period T =2*pi*r/V.. electric current is defined as the rate of flowing of charge. Hence I=Q/T......put the value of I=1*10^3, Q=5*10^-9,and T=2pi*r/V...and find V.

OpenStudy (anonymous):

thank you so much believe or not i couldnt solve this problem until now! thanks again